# Envelopes in the class of Banach algebras of polynomial growth and $C^{\infty}$-functions of a finite number of free variables

### v3

@@ 3. Envelopes of certain commutative algebras
Although our main object of study are envelopes of non-commutative algebras, we first consider the commutative case. The following result is the main reason that prompted the author to study envelopes with respect to $\mathsf{PG}$.
@@
A proof of a similar result for complex-valued functions of class $C^{\infty}$ and a complex Banach algebra can be found, for example, in [9, Theorem 1]. In the case of a real Banach algebra, it suffices to apply the result to the complexification of the algebra $B$ and note that $\Phi_{V}$ sends all the real-valued functions to $B$. The latter follows from the fact that $\mathbb{R}[x]$ is dense in $C^{\infty}(V)$. ∎
@@
It is obvious that Part (A) follows from Part (B). For a proof of Part (B) see [38, Chapitre VI, p. 119, Corollaire 1.5]. See also a proof of Part (A) in [28, Théorème 4] and partial cases in [22, 27].
@@ 4. Envelopes of certain non-commutative algebras
In general, the envelope of a commutative algebra with respect to $\mathsf{PG}$ can be treated as an algebra of $C^{\infty}$-functions on some commutative space and, moreover, the envelope of a non-commutative algebra as an algebra of $C^{\infty}$-functions on a non-commutative space. In this section, we consider first non-commutative examples: the envelopes of the coordinate algebras of quantum planes and quantum groups $SL_{q}(2,\mathbb{R})$ and also of the universal enveloping algebras of finite-dimensional Lie algebras. But we begin with another example.
@@
Taking the tensor product of the natural embeddings $\mathcal{R}(\Omega)\to C^{\infty}(\Omega)$ and $\mathbb{R}[u]\to\mathbb{R}[[u]]$ we obtain an injective linear map
@@
We now show that it can be treated as an enveloping homomorphism. Denote $C^{\infty}(\Omega)\mathbin{\widehat{\otimes}}\mathbb{R}[[u]]$ by $C^{\infty}(\mathbb{R}^{2}_{q})$ ($q\in\mathbb{R}\setminus\{0,1\}$).
@@
Denote $C^{\infty}(\mathbb{R}^{\times})\mathbin{\widehat{\otimes}}\mathbb{R}[[b,c]]$ by $C^{\infty}(SL_{q}(2,\mathbb{R}))$.
@@
(C) The algebra $C^{\infty}_{\mathfrak{g}}$ is independent of the choice of a basis in $\mathfrak{n}$ and its complement to a basis in $\mathfrak{g}$.
@@
(A) It is easy to see that $\mathfrak{g}^{\mathbf{tri}}=0$ when $\mathfrak{g}=\mathfrak{s}\mathfrak{l}_{2}$ or $\mathfrak{g}=\mathbb{R}^{2}\rtimes\mathfrak{s}\mathfrak{l}_{2}$. Thus $\widehat{U}(\mathfrak{g})^{\,\mathsf{PG}}$ is also trivial in both cases.
@@
Recall that $\mathcal{R}(SL_{q}(2,\mathbb{R}))$ and $U(\mathfrak{g})$ have Hopf algebra structures. Similarly, $\widehat{U}(\mathfrak{g})^{\,\mathsf{PG}}$ and ${\mathcal{R}(SL_{q}(2,\mathbb{R}))^{^}}{\,\,}^{\,\mathsf{PG}}$ admit induced topological Hopf algebra structures. This topic will be discuss in a separate paper.
@@ 5. Envelopes of free algebras
For a proof (in the case of a free Lie algebra over an arbitrary commutative ring) see [8, p. 64, § 2.4.2, Corollary 2.16(ii)] or [12, p. 72, Lemma 2.11.16].
@@
where $a$ and $b$ are elements of a Banach algebra and $s$ is a scalar; see, e.g., [10, Chapter II, § 15, p. 83, Remark 1].
@@
is absolutely integrable. Thus, the integral exists not only in the weak sense (of Gelfand and Pettis), but also in the strong sense (of Bochner) [21, p. 80, Theorem 3.7.4]. Applying Theorem 4.5 again, now to the tuple $(\mathop{\mathrm{ad}}\nolimits b_{1},\ldots,\mathop{\mathrm{ad}}\nolimits b_{k})$, we see that $\Psi_{\delta}$ is indeed an extension of the ordered calculus in (5.4). From the definition of $C^{\infty}(X)$ (see (5.5)), we have an extension of the map $V\to B$ to $C^{\infty}(X)\to B$.
@@
Let $m\geqslant 1$. Note that $\widetilde{\theta}_{\boldsymbol{\lambda},w}(a)=\theta_{w}(a)$ for every $a\in[\mathfrak{f}_{2},\mathfrak{f}_{2}]$ and so
$\widetilde{\theta}_{\boldsymbol{\lambda},w}(g_{\beta^{1}}\cdots g_{\beta^{m}})=\theta_{w}(g_{\beta^{1}})\cdots\theta_{w}(g_{\beta^{m}}).$
Note that the product of $m$ strictly upper-diagonal matrices have zero entries everywhere except possibly the upper-right corner. This entry depends only on the entries on the second diagonals and equals the product of the $p$th entries in the $p$th matrix, where $p$ runs $\{1,\ldots,m\}$. Then it follows from Lemma 5.8 that (5.17) holds. ∎
@@
By Lemma 5.9,
@@
It follows from a theorem of Dugundji [17, Theorem 4.1] that a continuous linear map between Fréchet spaces is topologically injective if and only if it has a continuous left inverse map (possibly, non-linear).

### v4

@@ 3. Envelopes of certain commutative algebras
Although our main object of study are envelopes of non-commutative algebras, we first consider the commutative case. The following result is the main reason that led the author to study envelopes with respect to $\mathsf{PG}$.
@@
A proof of a similar result for complex-valued functions of class $C^{\infty}$ and a complex Banach algebra can be found, for example, in [10, Theorem 1]. In the case of a real Banach algebra, it suffices to apply the result to the complexification of the algebra $B$ and note that $\Phi_{V}$ sends all the real-valued functions to $B$. The latter follows from the fact that $\mathbb{R}[x]$ is dense in $C^{\infty}(V)$. ∎
@@
It is obvious that Part (A) follows from Part (B). For a proof of Part (B) see [39, Chapitre VI, p. 119, Corollaire 1.5]. See also a proof of Part (A) in [29, Théorème 4] and partial cases in [23, 28].
@@ 4. Envelopes of certain non-commutative algebras
In general, the envelope of a commutative algebra with respect to $\mathsf{PG}$ can be treated as an algebra of $C^{\infty}$-functions on some commutative space and, moreover, the envelope of a non-commutative algebra as an algebra of $C^{\infty}$-functions on a non-commutative space. In this section, we consider first non-commutative examples: the envelopes of the coordinate algebras of quantum planes and quantum groups $SL_{q}(2,\mathbb{R})$ and also of the universal enveloping algebras of finite-dimensional Lie algebras. But we start with another example.
@@
Taking the tensor product of the natural embeddings $\mathcal{R}(\Omega)\to C^{\infty}(\Omega)$ and $\mathbb{R}[u]\to\mathbb{R}[[u]]$ (the latter denotes the space of formal power series), we obtain an injective linear map
@@
We now show that it can be treated as an enveloping homomorphism. Denote $C^{\infty}(\Omega)\mathbin{\widehat{\otimes}}\mathbb{R}[[u]]$ by $C^{\infty}(\mathbb{R}^{2}_{q})$ ($q\in\mathbb{R}\setminus\{0,1\}$). Note that $C^{\infty}(\mathbb{R}^{2}_{q})$ does not depend on $q$ as a locally convex space but the multiplication introduced in the following theorem does.
@@
Denote $C^{\infty}(\mathbb{R}^{\times})\mathbin{\widehat{\otimes}}\mathbb{R}[[b,c]]$ by $C^{\infty}(SL_{q}(2,\mathbb{R}))$. (To show that this space can be endowed with a multiplication that depends on $q$, we again use the index $q$.)
@@
(C) The algebra $C^{\infty}_{\mathfrak{g}}$ is independent of the choice of a basis in $[\mathfrak{g},\mathfrak{g}]$ and its complement to a basis in $\mathfrak{g}$.
@@
(A) It is easy to see that $\mathfrak{g}^{\mathbf{tri}}=0$ when $\mathfrak{g}=\mathfrak{s}\mathfrak{l}_{2}$ or $\mathfrak{g}=\mathbb{R}^{2}\rtimes\mathfrak{s}\mathfrak{l}_{2}$. Thus $\widehat{U}(\mathfrak{g})^{\,\mathsf{PG}}$ is also trivial in both cases. In general $\widehat{U}(\mathfrak{g})^{\,\mathsf{PG}}=0$ for every semi-simple real Lie algebra $\mathfrak{g}$.
@@
Recall that $\mathcal{R}(SL_{q}(2,\mathbb{R}))$ and $U(\mathfrak{g})$ have Hopf algebra structures. Similarly, $\widehat{U}(\mathfrak{g})^{\,\mathsf{PG}}$ and ${\mathcal{R}(SL_{q}(2,\mathbb{R}))^{^}}{\,\,}^{\,\mathsf{PG}}$ admit induced topological Hopf algebra structures. This topic is discussed in [8].
@@ 5. Envelopes of free algebras
For a proof (in the case of a free Lie algebra over an arbitrary commutative ring) see [9, p. 64, § 2.4.2, Corollary 2.16(ii)] or [13, p. 72, Lemma 2.11.16].
@@
where $a$ and $b$ are elements of a Banach algebra and $s$ is a scalar; see, e.g., [11, Chapter II, § 15, p. 83, Remark 1].
@@
is absolutely integrable. Thus, the integral exists not only in the weak sense (of Gelfand and Pettis), but also in the strong sense (of Bochner) [22, p. 80, Theorem 3.7.4]. Applying Theorem 4.5 again, now to the tuple $(\mathop{\mathrm{ad}}\nolimits b_{1},\ldots,\mathop{\mathrm{ad}}\nolimits b_{k})$, we see that $\Psi_{\delta}$ is indeed an extension of the ordered calculus in (5.4). From the definition of $C^{\infty}(X)$ (see (5.5)), we have an extension of the map $V\to B$ to $C^{\infty}(X)\to B$.
@@
Let $m\geqslant 1$. It is obvious that $\theta_{w}(g_{\beta^{1}}\cdots g_{\beta^{m}})=\theta_{w}(g_{\beta^{1}})\cdots\theta_{w}(g_{\beta^{m}})$. Note that the product of $m$ strictly upper-diagonal matrices have zero entries everywhere except possibly the upper-right corner. This entry depends only on the entries on the second diagonals and equals the product of the $p$th entries in the $p$th matrix, where $p$ runs $\{1,\ldots,m\}$. Then it follows from Lemma 5.8 that (5.17) holds. ∎
@@
Note also that $\widetilde{\theta}_{\boldsymbol{\lambda},w}(a)=\theta_{w}(a)$ for every $a\in[\mathfrak{f}_{2},\mathfrak{f}_{2}]$. Then by Lemma 5.9,
@@
It follows from a theorem of Dugundji [18, Theorem 4.1] that a continuous linear map between Fréchet spaces is topologically injective if and only if it has a continuous left inverse map (possibly, non-linear).